(x×4)(x+7)(x+8)(x+11)+20=0 Please solve above equation for x.

Grade:11

1 Answers

Arun
6 years ago
= 0Solving for variable 'x'.4 + x3 + 30x2 = 0Solving2484 + 1500x + 325x4 + x3 + 30x2 = 0Combine like terms: 2464 + 20 = 24842484 + 1500x + 325x4 + x3 + 30x2) + 20 = 0Reorder the terms:2464 + 20 + 1500x + 325x4 + x3 + 30x2(2464 + 1500x + 325x3 = 30x3 + 11x3) + 20 = 0Combine like terms: 19x4 + x3 + 11x3 + 19x2(2464 + 1500x + 325x2 = 325x2 + 209x2) + 20 = 0Combine like terms: 116x4 + x3 + 11x3 + 19x2 + 209x2)) + 20 = 0Combine like terms: 224x + 1276x = 1500x(2464 + 1500x + 116x4 + x3 + (11x3 + 19x2 + 209x2) * (11 + x)(2464 + 224x + 1276x + 116x3 + x2)(11 + x) + 20 = 0Multiply (224 + 116x + 19x3 + x2(224 + 116x + 19x2 = 19x2 + 4x2)(11 + x) + 20 = 0Combine like terms: 15x3 + x2 + 4x2))(11 + x) + 20 = 0Combine like terms: 56x + 60x = 116x(224 + 116x + 15x3 + x2 + (4x2) * (4 + x)(224 + 56x + 60x + 15x2)(4 + x)(11 + x) + 20 = 0Multiply (56 + 15x + x2))(4 + x)(11 + x) + 20 = 0Combine like terms: 8x + 7x = 15x(56 + 15x + x2Simplifying(x + 8)(x + 7)(x + 4)(x + 11) + 20 = 0Reorder the terms:(8 + x)(7 + x)(4 + x)(11 + x) + 20 = 0Multiply (8 + x) * (7 + x)(8(7 + x) + x(7 + x))(4 + x)(11 + x) + 20 = 0(56 + 8x + (7x + x we get there is no solution